# Forecasting and Manipulating the Forecasts of Others

### 18 Mar 2026 v2

@@ 2 The Decentralized LQG Game
The class of games we study has three features: (i) multiple agents share a common evolving state but observe it through private noisy channels; (ii) each agent’s action feeds back into the state dynamics and thereby into every other agent’s signal process; and (iii) agents minimize individual quadratic objectives over a finite horizon. Because actions enter the state, they simultaneously serve the physical objective and reshape what opponents observe, making each agent’s problem jointly strategic and informational. Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable—the same structure that underlies the workhorse models of [28, 26, 20] and the mean-field LQG literature [7], and that serves as the exact benchmark against which approximations can be assessed.
2.1 Primitives
@@
Observations.
Player $i$ does not observe $X_{t}$ directly. Instead, they observe a noisy measurement process $Y^{i}_{t}\in\mathbb{R}^{d}$ given by
$dY^{i}_{t}=\sqrt{P^{i}(t)}X_{t}\,dt+dW^{i}_{t},\quad Y^{i}_{0}=0,$ (2.2)
where: [^1] The square-root parameterization is convenient for interpreting $P^{i}(t)$ as a precision matrix, but all results extend to observation channels of the form $dY^{i}_{t}=H^{i}(t)X_{t}\,dt+E^{i}(t)\,dW_{t}$ with arbitrary deterministic $H^{i}(t)\in\mathbb{R}^{d_{i}\times d}$ and $E^{i}(t)E^{i}(t)^{\top}=I$, including correlated observation noises across players and signals that weight different state coordinates asymmetrically. All kernel formulas remain verbatim under this generalization.
  • $W^{i}_{t}\in\mathbb{R}^{d}$ is a standard Brownian motion (player $i$’s observation noise),
  • $P^{i}(t)\in\mathbb{R}^{d\times d}$ is symmetric positive semidefinite, and $\sqrt{P^{i}(t)}$ denotes its unique symmetric positive semidefinite square root.
  • All Brownian motions $\{W^{0},W^{1},\ldots,W^{n}\}$ are mutually independent.
Notation for primitive shocks.
We collect all primitive noises into a single vector $W_{t}:=(W^{0}_{t},\dots,W^{n}_{t})^{\top}\in\mathbb{R}^{(n+1)d}$. For each $j\in\{0,\dots,n\}$, the block selector $E^{j}\in\mathbb{R}^{d\times(n+1)d}$ extracts the $j$-th channel: $E^{j}W_{t}=W^{j}_{t}$. The block projector $\Pi^{i}:=E^{i,\top}E^{i}$ is the identity on the $i$-th block and zero elsewhere: $\Pi^{i}v=(0,\dots,0,v^{i},0,\dots,0)$. Player $i$ directly observes $\Pi^{i}W$ and must infer $(I-\Pi^{i})W$ from drift-based learning.
Unresolved impulse response.
For each player $i$, define
$\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u),$
the component of the state’s impulse response to the $u$-th shock that player $i$ has not yet resolved at time tt—large when the shock is recent or the observation channel weak, and shrinking as observations accumulate. Formally, $\widetilde{X}^{i}_{t}(u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 3.1), with its algebraic representation in terms of $F^{i}$ given in (A.12).
@@ 2.2 Nash Equilibrium
Throughout, a player’s choice is a strategy map: for each $t\in[0,T]$, the action $D_{t}^{i}$ is a (progressively measurable) functional of the private observation history $Y^{i}_{\cdot\wedge t}$. We write $D_{t}^{i}=D_{t}^{i}[Y^{i}]$, where the bracket notation $D_{t}^{i}[\cdot]$ denotes the map from the observation path $(Y^{i}_{s})_{s\leq t}$ to the action at time $t$. We identify a strategy with its induced adapted control process, assuming $\mathbb{E}\int_{0}^{T}\|D_{t}^{i}\|^{2}\,dt<\infty$.
@@
In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the noise-state linear class these maps are parameterized by deterministic maps that remain unchanged off equilibrium.
The requirement that opponents’ strategy maps be held fixed under deviations—rather than their realized actions or their realized beliefs—is the formal expression of the Lucas critique in this environment: if player $i$ changes their control law, the state dynamics change, which changes every opponent’s signal process, filtering problem, and actions, which feeds back into the state.
@@
The best-response closure (Corollary 3.5) does not require player $i$ to know opponents’ maps correctly: a player holding incorrect beliefs about opponents’ strategies still solves a well-posed best-response problem against the believed profile. A policy-maker can therefore let each agent optimize under its own beliefs, then evaluate welfare through the realized state dynamics (2.1), which depend on the actual strategy profile rather than on whether agents’ beliefs about each other are correct. Common knowledge of rationality is sufficient for Nash equilibrium but is not required to apply the characterization as a best-response operator.
@@
For each player $i$ and each $t\in[0,T]$, define the noise-state as the conditional mean path of the aggregated primitive noise:
@@
By construction, primitive-noise linear processes are $\mathbb{F}$-adapted Gaussian processes. Moreover, by Itô isometry the square-integrability condition above is equivalent to $L_{t}\in L^{2}$ for each $t$. Finally, the martingale representation theorem implies that any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable admits a representation of the form $\int_{0}^{t}Z_{s}\,dW_{s}$ with a stochastic integrand $Z$; the defining restriction of the linear class is that the response map $L(t,\cdot)$ is deterministic (for each fixed $t$).
@@
Fix a player $i$ and recall their noise-state path $\widehat{W}_{t}^{i}(\cdot)$ (Section 3). An admissible control $D^{i}=(D_{t}^{i})_{t\in[0,T]}$ is a noise-state linear control if it is affine in the noise state. That is, if there exist deterministic functions
@@
and we extend $D^{i}_{t}(u)$ by 0 for $u>t$. Equivalently, writing $\widehat{W}^{\,i}[Y^{i}]$ to emphasize dependence on the observation path, this defines the strategy map
@@ 3 Main results
We work in the linear–Gaussian model of Section 2. Two results drive the analysis. First, when opponents use noise-state linear strategies, the filtering equations inherit deterministic impulse responses (Theorem 3.1). Second, the best response over the full admissible class is itself noise-state linear (Theorem 3.2), so Nash equilibrium reduces to a deterministic fixed point in impulse-response maps. The filtering closure (Theorem 3.1) shows that noise-state dynamics have deterministic kernels; the best-response closure (Theorem 3.2) derives the adjoint system and the information wedge; Proposition 3.6 then decomposes equilibrium cost into a certainty-equivalent term and an information cost linked to endogenous precision.
@@
Let $\Pi^{i}:=E^{i,\top}E^{i}$ and $\widehat{X}_{t}^{i}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{i}]$, and define the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$.
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (3.2)
@@
$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du.$
@@
Discretizing time, each observation increment is affine in the noise increments $\Delta W_{1:k}$, so by joint Gaussianity the conditional mean $\mathbb{E}[\Delta W_{1:k}\mid\Delta Y^{i}_{1:k}]$ is a deterministic linear projection; each layer of “beliefs about beliefs” is a further deterministic composition of projections, which is why the regress closes. The decomposition in part (i) reflects a scale separation: the direct component $\Pi^{i}\,dW_{u}$ is $O(1)$ (player $i$ observes their own noise contemporaneously), while the indirect component—absolutely continuous in $u$ with density driven by $F^{i}$—captures drift-based inference about unobserved noise coordinates at scale $O(dt)$.
@@
Structure. Because the equilibrium control is a noise-state linear process $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,dW_{u}$ with deterministic mean and kernel, the backward system splits into two structurally identical subsystems—one for the mean coefficients $(\bar{H}^{X,i},\bar{H}^{k,i})$, one for the response-map coefficients $(H^{X,i},H^{k,i})$—that should be read in parallel.
@@
In each subsystem, the single-agent costate acquires an additional coupling term: player $i$’s action shifts the state, which shifts opponent $k$’s signal, posterior, and action, feeding back into the state. This channel introduces belief-adjoint variables $(\bar{H}^{k,i},H^{k,i})$ that track the shadow value of perturbing opponent $k$’s noise-state estimate, coupled into the physical costate through the opponents’ filtering gains.
@@
Mean backward system. The mean physical costate $\bar{H}^{X,i}_{t}$ and mean belief-adjoints $\bar{H}^{k,i}_{t}(u)$ satisfy
$\frac{d}{dt}\bar{H}^{X,i}_{t}$ $=-\big(G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\big)-A(t)^{\top}\bar{H}^{X,i}_{t}-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz}_{=:\;\bar{\mathcal{V}}^{i}_{t},\;\text{mean information wedge}},$ (3.3)
$\frac{d}{dt}\bar{H}^{k,i}_{t}(u)$ $=-\big(D^{k}_{t}(u)\big)^{\top}\bar{H}^{X,i}_{t}+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz,\qquad k\neq i,$ (3.4)
with terminal conditions $\bar{H}^{X,i}_{T}=G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}$ and $\bar{H}^{k,i}_{T}(\cdot)=0$.
@@
The first two terms in (3.3) are the standard single-agent costate equation; the third term $\bar{\mathcal{V}}^{i}_{t}$ is the mean information wedge, absent in the single-agent problem. In the belief-adjoint equation (3.4), the forcing $D^{k}_{t}(u)^{\top}\bar{H}^{X,i}_{t}$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}_{t}(u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock. The integral term is a self-coupling through which belief perturbations propagate forward via the filtering gain $\widetilde{X}^{k}$.
Kernel backward system. The kernel costate $H^{X,i}_{t}(r)$ and kernel belief-adjoints $H^{k,i}_{t}(u,r)$ satisfy the identical system with $\bar{X}_{t}\to X(t,r)$ and $\bar{G}^{X,i}_{t}\to G^{X,i}_{t}(r)$; the kernel information wedge $\mathcal{V}^{i}_{t}(r)$ is the corresponding substitution of $\bar{\mathcal{V}}^{i}_{t}$. We define the (full) information wedge as the linear process
$\mathcal{V}^{i}_{t}:=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r},$ (3.5)
which is the total distortion of the physical costate relative to the single-agent benchmark. The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ captures how belief manipulation distorts the equilibrium level (mean actions); the response-map wedge $\mathcal{V}^{i}_{t}(r)$ captures how it distorts the equilibrium response to shocks. Both components—and hence $\mathcal{V}^{i}_{t}$ itself—vanish when signals are exogenous.
Policy interpretation of the wedge.
The information wedge $\mathcal{V}^{i}_{t}$ vanishes identically when signals are exogenous to controls, recovering the standard LQG costate equation. When controls enter the state dynamics, a change in signal precision $P^{k}(t)$ also operates through the belief-manipulation channel: it alters how sharply opponents’ posteriors respond to drift perturbations (through $\widetilde{X}^{k}$), which alters the shadow value of shifting those posteriors (through $\bar{H}^{k,i}$), which feeds back into the equilibrium control through (3.3). Any model that omits this channel misattributes the strategic distortion to the physical response: the mean error is exactly $\bar{\mathcal{V}}^{i}_{t}$, and the shock-by-shock error is $\mathcal{V}^{i}_{t}(r)$. The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ governs the distortion in equilibrium mean actions (Section 4.3); the kernel wedge $\mathcal{V}^{i}_{t}(r)$ governs the distortion in equilibrium impulse responses. Section 4 quantifies both channels numerically.
@@
$G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u)=0,$
and hence
$D_{t}^{i}=-(G^{DD,i}(t))^{-1}\bar{H}^{X,i}_{t}-(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u).$
In particular, every best response over the full admissible class is itself a noise-state linear control.
@@
With opponents’ maps fixed and linear, the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion whose law does not depend on $D^{i}$ (Liptser–Shiryaev [16], Theorem 7.12). Since $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$ (Theorem 3.1 in innovations coordinates), player $i$’s control is a deterministic affine functional of the exogenous process $I^{i}$. The induced objective is strictly convex (Lemma 3.4), so the stationarity condition (B.8) is necessary and sufficient, yielding the stated formula. Note that the exogeneity of $I^{i}$ is a property of the best-response problem: the response-maps mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem C.5). ∎
@@
Since every best response over the full admissible $L^{2}$ class is noise-state linear (Corollary 3.5), any Nash equilibrium of the restricted game in which players are confined to noise-state linear controls is automatically a Nash equilibrium of the unrestricted game. The open question is whether equilibria outside the deterministic impulse-response class exist; Section 6 discusses a program for ruling them out.
@@
Risk-sensitive extension.
Replace (2.4) with the entropic risk measure $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$, which penalizes cost variance in addition to cost mean ($\theta_{i}>0$ governs risk aversion; $\theta_{i}\to 0$ recovers (2.4)). The entire framework carries over: the best-response FOC is identical to Corollary 3.5 with the noise-state $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ that tilts the conditional mean of $W$ toward directions that reduce expected cost and rescales conditional precision along directions where cost variance is large. The filtering closure is unchanged; certainty equivalence fails because the quadratic structure of the cost reshapes the posterior without altering the observation channel. The full statement and derivation are in Appendix B.6 (Theorem B.1).
Information cost and endogenous precision.
At the equilibrium profile, the control gap between full and partial information is
$D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$ (3.6)
where $D^{i,\mathrm{full}}_{t}:=-(G^{DD,i}(t))^{-1}M^{i}_{t}$ is the best response under full observation of $W$, $M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u}$ is the stochastic costate, and $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$ is the noise-state estimation error.
Proposition 3.6 (Information cost).
At the equilibrium profile,
$\Delta J^{i}:=J^{i}(D^{i};D^{*,-i})-J^{i}(D^{i,\mathrm{full}};D^{*,-i})=\int_{0}^{T}\mathrm{tr}\!\left[G^{DD,i}(t)^{-1}\,\Sigma^{M,i}_{t}\right]\mathrm{d}t,$ (3.7)
where $\Sigma^{M,i}_{t}:=\operatorname{Cov}(M^{i}_{t}\mid\mathcal{F}^{i}_{t})$ is the costate estimation error covariance (deterministic by Gaussianity). Moreover, $\Delta J^{i}$ depends on $P^{i}$ only through the filtering kernels $(\widetilde{X}^{i},F^{i})$ and is independent of player $i$’s control $D^{i}$.
The total cost decomposes as
$J^{i}(D^{i};D^{*,-i})=\underbrace{J^{i}(D^{i,\mathrm{full}};D^{*,-i})}_{\text{certainty-equivalent cost}}+\;\underbrace{\Delta J^{i}(P^{i})}_{\text{information cost}},$ (3.8)
This decomposition underlies the welfare comparisons in Section 4: the information wedge enters through $H^{X,i}$ inside $\Sigma^{M,i}_{t}$. The joint optimization over $(D^{i},P^{i})$ separates: $D^{i}$ is determined by Corollary 3.5 for any given $P^{i}$, and $P^{i}$ solves
$\min_{P^{i}(\cdot)\succeq 0}\quad\Delta J^{i}(P^{i})+\int_{0}^{T}c^{i}(t,P^{i}(t))\,\mathrm{d}t,$ (3.9)
where $c^{i}$ is a convex attention cost.
Corollary 3.7 (Endogenous precision closure).
If each opponent $k\neq i$ uses a deterministic precision path $P^{k}(\cdot)$, then $\Delta J^{i}$ is a deterministic functional of $P^{i}(\cdot)$, and the best-response precision is deterministic.

### 23 Mar 2026 v3

@@ 3 The Decentralized LQG Game
The games we study have three features: (i) multiple agents share a common evolving state but observe it through private noisy channels; (ii) each agent’s action feeds back into the state dynamics and thereby into every other agent’s signal process; and (iii) agents minimize individual quadratic objectives over a finite horizon. Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable.
3.1 State dynamics and objectives
@@
3.2 Information structure
Observations.
Player $i$ does not observe $X_{t}$ directly. Each player monitors the evolving state through private information sources: analyst reports, proprietary sensors, research output, market data feeds. Some sources arrive without cost (publicly posted prices, mandatory disclosures, direct observation of one’s own action) while others must be actively acquired or processed. The sources may be discrete (a finite set of news feeds) or continuous (a flow of independent measurements whose aggregate is naturally modeled by a Brownian sheet [17]).
In this paper we take each player’s information-processing capacity as exogenous: the menu of sources and the attention allocated to each are fixed and common knowledge. The observation process summarizing the player’s information about the state then takes the form
$dY^{i}_{t}=\Gamma^{i}(t)\,X_{t}\,dt+E^{i}\,dW_{t},\quad Y^{i}_{0}=0,$ (3.4)
where $\Gamma^{i}(t)\in\mathbb{R}^{d\times d}$ is a deterministic observation gain and $E^{i}\in\mathbb{R}^{d\times(n+1)d}$ is the block selector extracting player $i$’s noise channel ($W^{i}_{t}:=E^{i}W_{t}$). All Brownian motions $\{W^{0},W^{1},\ldots,W^{n}\}$ are mutually independent. The precision matrix $P^{i}(t):=\Gamma^{i}(t)^{\top}\Gamma^{i}(t)$ measures the Fisher information rate about the state per unit time; the gain $\Gamma^{i}$ is the primitive, and $P^{i}$ is derived. The block-selector form is without loss: replacing $E^{i}$ by a general invertible noise loading $\Sigma^{i}(t)$ amounts to substituting $\Gamma^{i,\top}(\Sigma^{i}\Sigma^{i,\top})^{-1}$ for $\Gamma^{i,\top}$ throughout.
Notation for primitive shocks.
We collect all primitive noises into a single vector $W_{t}:=(W^{0}_{t},\dots,W^{n}_{t})^{\top}\in\mathbb{R}^{(n+1)d}$. The block projector $\Pi^{i}:=E^{i,\top}E^{i}$ is the identity on the $i$-th block and zero elsewhere: $\Pi^{i}v=(0,\dots,0,v^{i},0,\dots,0)$. The block-selector structure ensures that each player’s observation noise is an independent component of $W$, giving the noise-state its direct/indirect decomposition (Theorem 4.1): player $i$ (almost) directly observes $\Pi^{i}W$ and infers $(I-\Pi^{i})W$ from drift-based learning.
Unresolved impulse response.
For each player $i$, define
$\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u),$
the component of the state’s impulse response to the $u$-th shock that player $i$ has not yet resolved at time tt—large when the shock is recent or the observation channel weak, and shrinking as observations accumulate. Formally, $\widetilde{X}^{i}_{t}(u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 4.1), with its explicit formula in (A.12).
Throughout Sections 45 the gain $\Gamma^{i}(\cdot)$ is fixed; Section 6 discusses the extension to endogenous precision.
@@ 3.3 Nash Equilibrium
Throughout, a player’s choice is a strategy map: for each $t\in[0,T]$, the action $D_{t}^{i}$ is a (progressively measurable) functional of the private observation history $Y^{i}_{\cdot\wedge t}$. We write $D_{t}^{i}=D_{t}^{i}[Y^{i}]$, where the bracket notation $D_{t}^{i}[\cdot]$ denotes the map from the observation path $(Y^{i}_{s})_{s\leq t}$ to the action at time $t$. We require $\mathbb{E}\int_{0}^{T}\|D_{t}^{i}\|^{2}\,dt<\infty$.
@@
In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the noise-state linear class these maps have deterministic coefficients that do not change when the state path changes.
Holding opponents’ strategy maps fixed under deviations (rather than their realized actions or beliefs) is substantive here because signals are endogenous: changing one player’s control law changes the state dynamics, which changes every opponent’s signal, filtering problem, and realized action, even though their decision rules have not changed.
@@
The best-response closure (Corollary 4.5) does not require player $i$ to know opponents’ maps correctly. A policy-maker can let each agent optimize under its own beliefs, then evaluate welfare through the realized state dynamics (3.1); common knowledge of rationality is sufficient for Nash equilibrium but not required to apply the characterization as a best-response operator.
@@
For each player $i$ and each $t\in[0,T]$, the noise-state encountered informally in Section 2.2 is defined as the conditional mean path of the aggregated primitive noise:
@@
By construction, primitive-noise linear processes are $\mathbb{F}$-adapted Gaussian processes. Moreover, by Itô isometry the square-integrability condition above is equivalent to $L_{t}\in L^{2}$ for each $t$. By the martingale representation theorem, any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable has a stochastic integral representation against $W$; the defining restriction of the linear class is that the integrand $L(t,\cdot)$ is deterministic. Equivalently, Definition 3.2 is the continuous-time analogue of an MA($\infty$) representation: $L(t,s)$ plays the role of the moving-average coefficient at lag tst-s, with the Itô integral replacing the infinite sum against i.i.d. innovations.
@@
Fix a player $i$ and recall their noise-state path $\widehat{W}_{t}^{i}(\cdot)$ (3.5). An admissible control $D^{i}=(D_{t}^{i})_{t\in[0,T]}$ is a noise-state linear control if it is affine in the noise state. That is, if there exist deterministic functions
@@
and we extend $D^{i}_{t}(u)$ by 0 for $u>t$. Equivalently, this defines the strategy map
@@ 4 Main results
This section derives the objects previewed in Section 2 in the general $n$-player model of Section 3.
The strategy is analogous to Harsanyi’s [12]: just as a common prior over a type space closes the belief hierarchy in static incomplete-information games, the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$ closes it here, with all higher-order conditional expectations following by linear projection onto deterministic impulse-response maps. The difference is that Harsanyi’s types are drawn once while noise unfolds continuously and actions reshape signals, so the closure requires a fixed point: the maps governing belief updating must be consistent with the strategy profile they induce (Figure 7).
@@
Let $\Pi^{i}:=E^{i,\top}E^{i}$ and $\widehat{X}_{t}^{i}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{i}]$, and define the innovation $dI_{t}^{i}:=dY_{t}^{i}-\Gamma^{i}(t)\,\widehat{X}_{t}^{i}\,dt$.
@@
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\;+\;E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (4.2)
@@
$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)\,dI_{t}^{i}\,du.$
@@
Discretizing time, each observation increment is affine in the noise increments $\Delta W_{1:k}$, so by joint Gaussianity the conditional mean $\mathbb{E}[\Delta W_{1:k}\mid\Delta Y^{i}_{1:k}]$ is a deterministic linear projection; each layer of “beliefs about beliefs” is a further deterministic composition of projections, which is why the regress closes. The decomposition in part (i) reflects a scale separation: the direct component $\Pi^{i}\,dW_{u}$ is $O(1)$ (player $i$ observes their own noise contemporaneously), while the indirect component captures drift-based inference about unobserved shocks at scale $O(dt)$.
@@
Structure. All objects below are linear processes (Definition 3.2): deterministic mean plus a stochastic integral against $W$ with deterministic kernel. The single-agent costate acquires an additional coupling term: player $i$’s action shifts the state, which shifts opponent $k$’s signal, posterior, and action, feeding back into the state. This channel introduces belief-adjoint processes $H^{k,i}_{t}(u)$ that track the shadow value of perturbing opponent $k$’s noise-state estimate.
@@
Backward system. The physical costate $H^{X,i}_{t}$ and belief-adjoints $H^{k,i}_{t}(u)$ are linear processes satisfying
$\frac{d}{dt}H^{X,i}_{t}$ $=-\big(G^{XX,i}(t)X_{t}+G^{X,i}_{t}\big)-A(t)^{\top}H^{X,i}_{t}-\underbrace{\sum_{k\neq i}P^{k}(t)\int_{0}^{t}\widetilde{X}^{k}_{t}(z)\,H^{k,i}_{t}(z)\,dz}_{=:\;\mathcal{V}^{i}_{t},\;\text{information wedge}},$ (4.3)
$\frac{d}{dt}H^{k,i}_{t}(u)$ $=-\big(D^{k}_{t}(u)\big)^{\top}H^{X,i}_{t}+X_{t}(u)^{\top}P^{k}(t)\int_{0}^{t}\widetilde{X}^{k}_{t}(z)\,H^{k,i}_{t}(z)\,dz,\qquad k\neq i,$ (4.4)
[…]
@@
The first two terms in (4.3) are the standard single-agent costate equation; the third term $\mathcal{V}^{i}_{t}$ is the information wedge, absent in the single-agent problem. In the belief-adjoint equation (4.4), the forcing $D^{k}_{t}(u)^{\top}H^{X,i}_{t}$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}_{t}(u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock. The integral term is a self-coupling through which belief perturbations propagate forward via the filtering gain $\widetilde{X}^{k}$; the same inner product $\int_{0}^{t}\widetilde{X}^{k}_{t}(z)\,H^{k,i}_{t}(z)\,dz$ that defines the wedge in (4.3) drives the self-coupling in (4.4), so the belief-adjoint feeds back into itself through the channel it prices.
Since all coefficients in (4.3)–(4.4) are deterministic, taking expectations recovers a closed system for the mean coefficients $(\bar{H}^{X,i},\bar{H}^{k,i})$; matching the coefficient of $dW_{r}$ recovers a closed system for the response-map coefficients $(H^{X,i}(\cdot,r),H^{k,i}(\cdot,\cdot,r))$. The two subsystems decouple and are solved separately in the numerical scheme of Section 2. The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ is the general-$n$ counterpart of the object displayed in Figures 45; the kernel wedge $\mathcal{V}^{i}_{t}(r)$ distorts equilibrium impulse responses. Both vanish when signals are exogenous.
@@
Remark 4.2 (Mean-field approximation error).
Mean-field models zero both wedge components by construction: in the continuum limit no single agent’s action measurably affects any other agent’s signal. Eliminating both components in a finite-$n$ model requires either exogenous signals (Corollary 4.3) or $n\to\infty$ with vanishing individual impact—the first is ruled out when information is costly [10], and concentration of informed trading pushes real markets toward finite $n$. In the Kyle–Back embedding (Section 5), the sign symmetry $W\mapsto-W$ zeros the mean wedge while the kernel wedge remains active; this symmetry breaks in any market with asymmetric positions.
@@
$G^{DD,i}(t)D_{t}^{i}+\mathbb{E}[H^{X,i}_{t}\mid\mathcal{F}_{t}^{i}]=0,\qquad\text{i.e.,}\qquad D_{t}^{i}=-(G^{DD,i}(t))^{-1}\mathbb{E}[H^{X,i}_{t}\mid\mathcal{F}_{t}^{i}].$
Since $H^{X,i}_{t}$ is a linear process and $\mathbb{E}[H^{X,i}_{t}\mid\mathcal{F}_{t}^{i}]=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u)$, every best response over the full admissible class is itself a noise-state linear control.
@@
With opponents’ maps fixed and linear, the innovation $dI_{t}^{i}:=dY_{t}^{i}-\Gamma^{i}(t)\,\widehat{X}_{t}^{i}\,dt$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion whose law does not depend on $D^{i}$ (Liptser–Shiryaev [19], Theorem 7.12). Since $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$ (Theorem 4.1 in innovations coordinates), player $i$’s control is a deterministic affine functional of the exogenous process $I^{i}$. The induced objective is strictly convex (Lemma 4.4), so the stationarity condition (B.8) is necessary and sufficient, yielding the stated formula. The exogeneity of $I^{i}$ is a property of the best-response problem: the response-maps mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem D.5). ∎
@@
Since every best response over the full admissible $L^{2}$ class is noise-state linear (Corollary 4.5), any Nash equilibrium in noise-state linear controls is automatically a Nash equilibrium of the unrestricted game; whether equilibria outside this class exist is discussed in Section 6.
Remark 4.4 (Single-player reduction).
@@
Proposition 4.6 (Failure of separation).
For $n\geq 2$ with endogenous signals, the equilibrium policy kernel $D^{i}_{t}(u)$ does not factor as $S(t)\cdot X_{t}(u)$ for any matrix-valued function $S$, unless the information wedge vanishes (Corollary 4.3). A change in signal precision alters the policy kernel itself, not merely the estimate to which it is applied.